Shift report 20000718 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Tue Jul 18 2000 - 07:45:00 EDT

  • Next message: Konstantin Olchanski: "[pisarski@quark.phy.bnl.gov: phobos results]"

    ------------------------------------------------------------
    Date: Tue Jul 18 07:45:00 2000, Shift: 00:00-08:00
    Supervisor: BS, with: TV
    Log book#3 pages 25 - 26
    
    Shift Summary:
    ------------------------------------------------------------
    A very slow shift... No beam at all, except a few short
    injections. RHIC practiced fast ramping - a lot faster
    than before - but they keep loosing the beam every
    time they start to ramp.
    
    Run 1805 is finished with 10000 ZDC-pulser events.
    
    There were no injection-periods long enough to test any
    of the new detector-developments...
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Tue Jul 18 2000 - 07:54:54 EDT